OpenStudy (anonymous):

A small beach ball has a radius of 8 inches. A larger beach ball has a volume that is twice the volume of the small ball. What is the best approximation for the radius of the large beach ball?

OpenStudy (yanasidlinskiy):

Hello Mikaelab! Let x = the radius of the large ball. \[\frac{ 4 }{ 3 }\times \pi r^2=2(\frac{ 4 }{ 3 }\times \pi 8^3)\] Multiply both sides by 3 to get, to get rid of the denominator, Find 8^3 \[4(\pi \times r ^3) = 8(\pi \times 512)\] Divide both sides by \[4 \times \pi\] \[r^3=2(512)\] \[r^3=1024\] \[r = 3\sqrt{1024}\] r = 10.08 inches, radius of the larger.